OpenStudy (anonymous):

Identify the x and y intercepts for the equation 4x - 3y = 12. (4, 0) and (0, -3) (3, 0) and (0, -4) (4, 0) and (0, 3) (-3, 0) and (0, 4) (3, 0) and (0, 4)

OpenStudy (anonymous):

x-intercept: y = 0 4x - 3(0) = 12 4x = 12 x =12/4 = 3 y-intercept: x = 0 4(0) - 3y = 12 -3y = 12 y = 12/(-3) = -4 so, the coordinates of x-intercept and y-intercept are: (3, 0) and (0, -4)